OpenStudy (anonymous):

What is the common ratio of the following geometric sequence? 7/12, -1/4, 3/28, 9/196, 27/1327 A. -1/4 B. 1/4 C. 3/7 D. -3/7

OpenStudy (gorv):

divide 2nd term by first

OpenStudy (anonymous):

I don't understand how to do it. I'm not good in math

OpenStudy (gorv):

\[\frac{ -1 }{ 4 }\div \frac{ 7 }{ 12}=\frac{ -1 }{ 4 }*\frac{ 12 }{ 7}=\frac{ -3 }{ 7 }\]

OpenStudy (mathmale):

gorv's thinking is that he needs to find the common ratio. In other words, what is the number by which we multiply the first term to get the second? What happens if we multiply the 2nd term by this same ratio? Pretty please, Angela, don't demean yourself with negative remarks. You can learn this material!

OpenStudy (anonymous):

I really have no idea. Does it mean if you multiply it the outcome of both of them will be the same ?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

I got it thank you so much to both of you!